OpenStudy (anonymous):

Simplify: 9h – 12h = 54 – 23

OpenStudy (anonymous):

combine the h terms and then get the number attached to h to the right hand side

OpenStudy (anonymous):

and what is the h term again im dialectics and very bad with numbers and im just getting more and more confused attach what to what ???

OpenStudy (anonymous):

@PhysicsGuru

OpenStudy (solomonzelman):

Combine like terms, and divide both sides by '-3'.

OpenStudy (anonymous):

the terms that have h which is 9 and -12

OpenStudy (anonymous):

you mean 9h and 12h @SolomonZelman

OpenStudy (solomonzelman):

yes. and then 54 and – 23

OpenStudy (anonymous):

you substract 9h with 12h, and substract 54 with 23. after such, you should have 3h=31. You divide both sides by 3. It should give you h=31/3

OpenStudy (anonymous):

yes perform the operation in between the numbers

OpenStudy (anonymous):

@Dread dont give the answers directly please it defeats the purpose of learning

OpenStudy (anonymous):

Sorry/

OpenStudy (anonymous):

thank you @Dread even tho your answer is not an option and thank you @SolomonZelman @PhysicsGuru

OpenStudy (solomonzelman):

Anytime, as long as you are clear and comfortable with this problem, @hardcandi1.

OpenStudy (anonymous):

-3h = -31 is what im getting instead of h=-31/3

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

h=-31/3 is right